### v1

@@ Abstract
Threshold-Computation-in-the-Head (TCitH) and VOLE-in-the-Head (VOLEitH), two recent developments of the MPC-in-the-Head (MPCitH) paradigm, have significantly improved the performance of digital signature schemes in this framework.
@@
In this note, we embed the restricted decoding problem, denoted by $\mathcal{E}$-SDP, within these frameworks. We propose a structurally simple modeling that achieves competitive signature sizes. Specifically, by instantiating the restricted decoding problem with the same hardness assumption underlying CROSS, we reduce sizes by more than a factor of two compared to the NIST submission. Moreover, we observe that ternary full-weight decoding, closely related to the hardness assumption underlying WAVE, is a restricted decoding problem. Using ternary full-weight decoding, we obtain signature sizes comparable to the smallest MPCitH-based candidates in the NIST competition.
[…]
1 Introduction
@@
Following the standardization of the signatures CRYSTALS-Dilithium, FALCON, and SPHINCS+, NIST issued a call for additional digital signature schemes in September 2022. This initiative, to which 40 proposals were submitted, has the intention of diversifying the portfolio of post-quantum signature standards. In October 2024, NIST announced that fourteen submissions are going to proceed to round two. Among these, six are based on the Threshold Computation in-the-Head (TCitH) and the VOLE in-the-head (VOLEitH) frameworks [1, 2, 3, 5, 12, 14]. These recent specializations of the Multi-Party Computation (MPC) in-the-Head (MPCitH) paradigm offer significant performance improvements compared to previous constructions.
1.0.1 MPCitH.
@@
Originally proposed in 2007 by Ishai, Kushilevitz, Ostrovsky, and Sahai [23], the MPCitH techniques have reshaped the landscape of post-quantum digital signatures. In a nutshell, the MPCitH approach constructs digital signatures by combining secure multi-party computation (MPC) with the Fiat–Shamir transform. Emulating an MPC protocol, the prover (i.e., the signer) performs computations on shares of a secret. A zero-knowledge proof of knowledge is obtained by opening a verifier-selected subset of the parties, which is checked for consistency.
1.0.2 TCitH and VOLEitH.
@@
The TCitH framework [21] instantiates the MPCitH paradigm with an $\ell$-out-of-$n$ threshold secret sharing. Alternatively, the TCitH framework may be described in the Polynomial Interactive Oracle Proofs (PIOP) formalism, as presented in [20]. This formalism also encompasses the VOLEitH framework, originally introduced as a technique to compile VOLE-based zero-knowledge protocols into publicly verifiable protocols, from which signatures can be obtained via the Fiat–Shamir transform [13]. Using the TCitH and VOLEitH frameworks, performant signatures can be obtained from various hardness assumptions, in particular from coding theory. Submissions to the NIST competition utilize Hamming-metric [3] and rank-metric decoding problems [2, 5], as well as the permuted kernel problem [1].
1.0.3 Our Contribution.
@@
In this work, we consider the restricted decoding problem, called $\mathcal{E}$-SDP, underlying the CROSS signature scheme. We present a modeling of $\mathcal{E}$-SDP that allows constructing signatures in the TCitH and VOLEitH frameworks. While structurally simple, the proposed modeling achieves competitive performance. In fact, for the exact same decoding problem underlying CROSS, we reduce signature size by more than a factor of two. This shows that, by replacing the simple CVE-based proof of knowledge of CROSS with recent variants of the MPCitH framework, signature sizes can be reduced, albeit at the cost of a significantly more complex construction. Further, we consider the problem of ternary full-weight decoding, a special case of the hardness assumption underlying the WAVE signature scheme [10, 19]. Here, the proposed modeling achieves signature sizes that are comparable to the smallest ones of the NIST competition.
1.0.4 Organization.
Section 2.0.4 provides a concise review of the TCitH and VOLEitH frameworks. Table 3 presents the proposed modeling of $\mathcal{E}$-SDP, and discusses the achievable signature sizes.
1.0.5 Artifacts.
The accompanying repository [16] contains the code to reproduce the provided parameters.
2 Signatures from Threshold Computation in the Head
The VOLE-in-the-Head (VOLEitH) framework [13] and the Threshold-Computation-in-the-Head (TCitH) framework [21] are recent specializations of the MPC-in-the-Head paradigm [23]. Since we build our contribution in these frameworks, we provide a concise overview following the Polynomial Interactive Oracle Proof (PIOP) formalism presented in [20].
2.0.1 Notation.
For an integer $m\in\mathbb{Z}$, we denote $m]=\{1,\ldots,m\}$. Let $\mathbb{F}$ be the finite field with $p$ elements, where $p$ is a prime number, and $\mathbb{K}$ its degree-$\mu$ field extension. We denote by $\mathbb{F}[x_{1},\ldots,x_{n}]_{\leq d}$ the set of polynomials in variables $\mathbf{x}=(x_{1},\ldots,x_{n})$ with coefficients in $\mathbb{F}$ and total degree at most $d$. Vectors are denoted with bold lowercase letters, and matrices with bold uppercase letters. Polynomials are denoted in plain lowercase letters, and vectors of polynomials in plain uppercase letters. For a vector of polynomials $P(X)\in\mathbb{F}[X]_{\leq d}^{n}$, we denote by $P(\infty)\in\mathbb{F}^{n}$ the vector of leading (degree-$d$) coefficients.
2.0.2 TCitH in a Nutshell.
The general zero-knowledge proof of knowledge obtained from the TCitH framework is provided in Figure 1.
@@ 2 Signatures from Threshold Computation in the Head
The prover attempts to convince the verifier that they know a witness $\mathbf{w}\in\mathbb{F}^{n}$ satisfying $F\left(\mathbf{w}\right)=\mathbf{0}\in\mathbb{F}^{m}$ for a One-Way Function (OWF) $F$, described by the system of polynomial relations $F=(f_{1},\ldots,f_{m})\in\big(\mathbb{F}\left[x_{1},\ldots,x_{n}\right]_{\leq d}\big)^{m}$. To this end, the prover uses a GGM-based Vector Commitment (VC) scheme [11, 13] to sample vectors of random polynomials
$P_{u}\in\big(\mathbb{K}\left[X\right]_{\leq d-1}\big)^{\eta}\quad\text{and}\quad P_{w}\in\big(\mathbb{K}\left[X\right]_{\leq 1}\big)^{n}\quad\text{with}\quad P_{w}\left(\infty\right)\in\mathbb{F}^{n},$
@@
which also generates a commitment $\mathsf{com}$ to $P_{w},P_{u}$ and the opening key $\mathsf{key}$. To embed the witness into $P_{w}$, the prover computes the offset $\boldsymbol{\Delta}_{\mathbf{w}}=\mathbf{w}-P_{w}\left(\infty\right)\in\mathbb{F}^{n}$, for which $\tilde{P}_{w}\left(X\right)=P_{w}\left(X\right)+\boldsymbol{\Delta}_{\mathbf{w}}\cdot X$ has leading [^1] Embedding the witness into the constant coefficient is also possible; see [14]. coefficients $\mathbf{w}$. The offset $\boldsymbol{\Delta}_{\mathbf{w}}$ and commitment $\mathsf{com}$ are then sent to the verifier. Next, the polynomial relation $F$ is evaluated at $\tilde{P}_{w}\left(X\right)$. Specifically, for $i\in[m]$ and $\ell\leq d$, denote by $f_{i,\ell}(\mathbf{x})$ the degree-$\ell$ homogeneous component of $f_{i}(\mathbf{x})$. Then, $F(\tilde{P}_{w})(X)$ is defined as
$F(\tilde{P}_{w})(X)\coloneqq\left(\sum_{\ell}X^{d-\ell}f_{1,\ell}\left(\tilde{P}_{w}\left(X\right)\right),\ldots,\sum_{\ell}X^{d-\ell}f_{m,\ell}\left(\tilde{P}_{w}\left(X\right)\right)\right).$
@@
Then $F(\tilde{P}_{w})(\infty)=F\left(\mathbf{w}\right)$, i.e., the embedded witness satisfies the polynomial relation $F$ if and only if $F(\tilde{P}_{w})(X)$ has degree at most d1d-1. Hence, $F(\tilde{P}_{w})(X)$ consists of $m$ polynomials of degree d1d-1. The number of polynomials is reduced through batching. To this end, the verifier samples a random matrix $\boldsymbol{\Gamma}\in\mathbb{K}^{\eta\times m}$ and the prover computes $\boldsymbol{\Gamma}\cdot F(\tilde{P}_{w})$, thereby reducing the number of polynomials from $m$ to $\eta$ while preserving the degree with overwhelming probability. The verifier masks $\boldsymbol{\Gamma}\cdot F(\tilde{P}_{w})(X)$ by computing
$Q(X)=\boldsymbol{\Gamma}\cdot F(\tilde{P}_{w})(X)+P_{u}\left(X\right),$
@@
where $P_{u}\left(X\right)$ is the length-$\eta$ vector of random polynomials of degree (d1)(d-1) generated by the VC scheme. Sending the masked polynomial vector $Q(X)$ to the verifier does not reveal any information about the witness. To verify correctness, the verifier challenges the prover at a random evaluation point $r$ in the evaluation domain $\Omega\subseteq\mathbb{K}$ of size $N=\left\lvert\Omega\right\rvert$. The prover produces the corresponding opening $\mathsf{opening}_{r}$ using its opening key $\mathsf{key}$. The verifier checks $\mathsf{opening}_{r}$ against $\mathsf{com}$ and, in case of success, computes the evaluations $P_{w}\left(r\right)\in\mathbb{K}^{\eta}$ and $P_{u}\left(r\right)\in\mathbb{K}^{n}$. With these evaluations, the verifier computes $P_{u}\left(r\right)+\boldsymbol{\Gamma}\cdot F(\tilde{P}_{w}\left(r\right))$ and checks whether it matches $Q\left(r\right)$, where $Q(X)$ is the polynomial vector received from the prover.
Remark 1
@@
In [14], also a three-pass variant of Figure 1 is given, which is obtained by skipping the batching step. This simplifies the protocol, but increases the required communication, which is why we omit this variant.
2.0.3 TCitH signature scheme.
@@
The proof system above has a soundness error given by
@@
Property 1
For grinding parameter $w$, let the proof system parameters satisfy
@@
Then the resulting signature scheme is EUF-CMA-secure [14]. Signatures are of the form $\sigma=(\sigma_{\mathrm{sym}},\sigma_{\mathbf{w}},\sigma_{F})$, where $\sigma_{\mathbf{w}}$ corresponds to the witness, $\sigma_{F}$ to the OWF $F$, and $\sigma_{\mathrm{sym}}$ to the VC scheme. The components’ sizes [^2] In general, the sizes depend on the specific instantiation of the VC scheme [22], and optimizations not described here can be applied. The formulas given follow [14], to which we refer the reader for further details. are given by
$\left\lvert\sigma_{\mathrm{sym}}\right\rvert$ $=\tau\cdot\lambda\cdot(\log_{2}N+1)+5\lambda+32,$
$\left\lvert\sigma_{\mathbf{w}}\right\rvert$ $=\tau\cdot n\cdot\log_{2}p,$
$\left\lvert\sigma_{F}\right\rvert$ $=\tau\cdot\eta\cdot(d-1)\cdot\mu\cdot\log_{2}p.$
2.0.4 VOLEitH.
The VOLEitH framework is closely related to the TCitH framework described above. Instead of performing $\tau$ parallel repetitions, VOLEitH employs a technique that allows sampling $r$ from an evaluation domain $\Omega$ of size
@@
Avoiding parallel repetitions reduces the soundness error, which in turn enables shorter signatures. However, this comes at the cost of a consistency check, an additional round of interaction between the prover and the verifier. For further details, we refer to [12, 13].
@@
$\frac{N^{\tau}}{d}\geq 2^{\lambda-w}\quad\text{and}\quad p^{\rho}\geq 2^{\lambda},\quad\text{with}\quad\rho=\tau\cdot\left\lceil\log_{p}N\right\rceil,\quad B=\left\lceil\frac{16}{\log_{2}p}\right\rceil.$
@@
$\left\lvert\sigma_{\mathrm{sym}}\right\rvert$ $=\tau\cdot(\rho+B)\log_{2}p+\tau\cdot 2\lambda+T_{\mathrm{open}}\cdot\lambda+4\lambda+32,$
$\left\lvert\sigma_{\mathbf{w}}\right\rvert$ $=\tau\cdot n\cdot\log_{2}p,$
$\left\lvert\sigma_{F}\right\rvert$ $=\tau\cdot(d-1)\cdot\rho\cdot\log_{2}p.$
Similar to the TCitH framework, which requires $\eta\cdot\mu\cdot\log_{2}p\geq\lambda$, the VOLEitH framework requires $\rho\cdot\log_{2}p\geq\lambda$. However, VOLEitH allows for a smaller $\tau$, owing to its reduced soundness error, arguably at the cost of a more complex scheme.
3 Modeling Restricted Decoding Problems
In this section, we present a so-called modeling for the restricted decoding problem, i.e., we express the OWF as a polynomial relation that can be used in the TCitH and VOLEitH frameworks. Before doing so, we briefly recap some basic notions from coding theory.
3.0.1 Coding theory.
A linear code over $\mathbb{F}$ of length $n$ is a $k$-dimensional subspace of $\mathbb{F}^{n}$. The code can be represented by a parity-check matrix $\mathbf{H}\in\mathbb{F}^{(n-k)\times n}$ such that $\mathbf{c}\mathbf{H}^{\top}=\mathbf{0}$ for every codeword $\mathbf{c}$. We denote by r=nkr=n-k the redundancy of the code. The syndrome of a vector $\mathbf{v}$ (through $\mathbf{H}$) is $\mathbf{s}=\mathbf{v}\mathbf{H}^{\top}\in\mathbb{F}^{r}$. The Hamming weight (from now on, only weight) of a vector $\mathbf{v}$ is the number of non-zero entries in $\mathbf{v}$.
3.0.2 Restricted decoding assumption.
The (general) restricted Syndrome Decoding Problem ($\mathcal{E}$-SDP) [8] replaces the Hamming-weight constraint of the classical SDP with the requirement that each entry must lie in the fixed subset $\mathcal{E}$ of the underlying finite field $\mathbb{F}$. Formally, it is defined as follows:
Definition 1($\mathcal{E}$-SDP)
Let $\mathcal{E}\subset\mathbb{F}$ denote the restriction, $\mathbf{H}\in\mathbb{F}^{r\times n}$ the parity-check matrix, and $\mathbf{s}\in\mathbb{F}^{r}$ the syndrome. The $\mathcal{E}$-SDP instance $(\mathbf{H},\mathbf{s})$ is solved by $\mathbf{e}\in\mathcal{E}^{n}$ with $\mathbf{e}\mathbf{H}^{\top}=\mathbf{s}$.
We refer to the size of the restriction as $z=\left\lvert\mathcal{E}\right\rvert$. Different choices of $\mathcal{E}$ are possible, and all of them lead to an NP-complete problem [9]. Restrictions of the form $\mathcal{E}=\{-a,\ldots,a\}$ for $a\ll p$ yield a problem similar in nature to learning with errors (LWE) [8], while [9] considers cyclic subgroups of $\mathbb{F}^{*}=\mathbb{F}\setminus\{0\}$. In general, any choice of restriction requires a dedicated analysis of the problem’s concrete computational hardness. Therefore, we focus on specific choices that have undergone careful analysis.
CROSS restriction.
The signature scheme CROSS [6] uses the restriction $\mathcal{E}=\{2^{i}\mid i\in[7]\}$ within $\mathbb{F}$ of size $p=127$. We are going to refer to this setting as CROSS-SDP, the concrete hardness of which is analyzed in detail in [7] and confirmed in [15]. For instance, the parameters $n=127$ and $k=76$ provide a suitable choice for NIST security category 1.
Ternary decoding with full weight.
The signature scheme WAVE [19, 10] relies on the ternary syndrome decoding with high weight. For state-of-the-art solvers [17], parameterizing the problem such that the error is full-weight and unique, maximizes the concrete hardness (a setting in which WAVE cannot sign successfully). This corresponds to $\mathcal{E}$-SDP with $\mathcal{E}=\{1,2\}$ and $p=3$, to which we are going to refer as Ternary-SDP. For $k\approx 0.369\cdot n$, the solver presented in [17] requires $\mathcal{O}\left(2^{0.247\cdot n}\right)$ operations, up to a polynomial factor. In particular, the parameters $n=579$ and $k=213$ are shown to achieve NIST security category 1.
3.0.3 Modeling $\mathcal{E}$-SDP.
The linearity of the syndrome equations allows a standard optimization that reduces the witness size from $n$ to $k$. We assume that the parity-check matrix is in systematic form [^4] This assumption, also made e.g. in [4], does not impact the hardness of $\mathcal{E}$-SDP. $\mathbf{H}=(\mathbf{A},\mathbf{1}_{r})$, where $\mathbf{A}\in\mathbb{F}^{r\times k}$ and $\mathbf{1}_{r}\in\mathbb{F}^{r\times r}$ denotes the identity matrix. Then, $\mathbf{e}\in\mathbb{F}^{n}$ satisfies $\mathbf{e}\mathbf{H}^{\top}=\mathbf{s}$ if and only if it is of the form $\mathbf{e}=(\mathbf{w},\mathbf{s}-\mathbf{w}\mathbf{A}^{\top})$ with $\mathbf{w}\in\mathbb{F}^{k}$. That is, the witness is defined as a length-$k$ partial error vector that is expanded to a length-$n$ error vector so that the syndrome equations are satisfied.

### v2

@@ Abstract
Threshold-Computation-in-the-Head (TCitH) and VOLE-in-the-Head (VOLEitH), two recent developments of the MPC-in-the-Head (MPCitH) paradigm, have significantly improved the performance of digital signature schemes. This work embeds the restricted decoding problem within these frameworks: we propose a structurally simple modeling that achieves competitive signature sizes. Specifically, by instantiating the restricted decoding problem with the same hardness assumption underlying CROSS, we reduce sizes by more than a factor of two compared to the NIST submission. Moreover, we observe that ternary full-weight decoding, closely related to the hardness assumption underlying WAVE, is a restricted decoding problem. Using ternary full-weight decoding, we obtain signature sizes comparable to the smallest MPCitH-based candidates in the NIST competition.
I Introduction
@@
Post-quantum signatures. Following the standardization of the signatures Dilithium (ML-DSA), FALCON (FN-DSA), and SPHINCS+ (SLH-DSA), NIST issued a call for additional digital signature schemes in September 2022. In October 2024, NIST announced that fourteen submissions proceeded to the second round. Among these, six are based on the Threshold Computation in-the-Head (TCitH) and the Vector Oblivious Linear Evaluation (VOLE) in-the-head (VOLEitH) technique [1, 2, 3, 4, 5, 6]. These recent specializations of the Multi-Party Computation (MPC) in-the-Head (MPCitH) paradigm offer significant performance improvements.
@@
MPCitH. Originally proposed in 2007 by Ishai, Kushilevitz, Ostrovsky, and Sahai [7], the MPCitH techniques have reshaped the landscape of post-quantum digital signatures. In a nutshell, the MPCitH approach constructs digital signatures by combining secure MPC with the Fiat–Shamir transform. Emulating an MPC protocol, the prover (i.e., the signer) performs computations on shares of a secret. A zero-knowledge proof of knowledge is obtained by opening a verifier-selected subset of the parties, which is checked for consistency.
@@
TCitH and VOLEitH. The TCitH framework [8] instantiates the MPCitH paradigm with an $\ell$-out-of-$n$ threshold secret sharing. Alternatively, the TCitH framework may be described in the Polynomial Interactive Oracle Proofs (PIOP) formalism, as presented in [9]. This formalism also encompasses the VOLEitH framework, originally introduced as a technique to compile VOLE-based zero-knowledge protocols into publicly verifiable protocols, from which signatures can be obtained via the Fiat–Shamir transform [10].
Code-based Cryptography refers to a class of cryptographic protocol whose security relies on hard problems in coding theory, such as syndrome decoding or code equivalence. In the context of the ongoing NIST competition, six out of the fourteen submissions are based on different coding-theory related assumption, such as code equivalence [11], restricted decoding [12], Hamming-metric decoding [3], rank-metric decoding decoding [2, 4], and the permuted kernel problem [1]. Among these, the last four are based on either TCitH or VOLEitH.
@@
Our Contribution. In this work, we consider the restricted decoding problem, called $\mathcal{E}$-SDP, underlying the CROSS signature scheme [12]. We present a modeling of $\mathcal{E}$-SDP that allows constructing signatures in the TCitH and VOLEitH frameworks. By doing so, we reduce CROSS signature sizes by more than a factor of two, at the cost of a more complex construction. Further, we consider the problem of ternary full-weight decoding, a special case of the hardness assumption underlying the WAVE signature scheme [13, 14]. Here, the proposed modeling achieves signature sizes comparable to the smallest [^1] Except for SQISign [15], which is considerably shorter and slower. Zero-Knowledge-based (ZK-based) schemes of the NIST competition.
Organization. Section˜II provides a review of the TCitH and VOLEitH frameworks. Section˜III presents the proposed modeling of $\mathcal{E}$-SDP, and Section˜IV provides a detailed performance evaluation. The repository [16] contains code to reproduce the provided figures and parameters.
II TCitH and VOLEitH frameworks
@@
We provide a concise overview following the PIOP formalism presented in [9] used to build TCitH and VOLEitH signatures.
Notation. $\lambda$ denotes the security parameter; NIST (security) categories 1, 3, and 5 correspond to $\lambda=143$, $\lambda=207$, and $\lambda=272$, respectively. For an integer $m$, we denote by $m]=\{1,\ldots,m\}$. Let $\mathbb{F}_{p}$ be the finite field with $p$ elements, where $p$ is a prime number, and $\mathbb{K}$ its degree-$\mu$ field extension. We denote by $\mathbb{F}_{p}[x_{1},\ldots,x_{k}]_{\leq d}$ the set of polynomials in variables $\mathbf{x}=(x_{1},\ldots,x_{k})$ with coefficients in $\mathbb{F}_{p}$ and total degree at most $d$. Vectors are denoted with bold lowercase letters, and matrices with bold uppercase letters. Polynomials are denoted in plain lowercase letters, and vectors of polynomials in plain uppercase letters. For a vector of polynomials in the variable $X$ $P(X)\in\mathbb{F}_{p}[X]_{\leq d}^{k}$, we denote by $P(\infty)\in\mathbb{F}_{p}^{k}$ the vector of leading (degree-$d$) coefficients. We write $x\propto y$ to denote that $x$ is proportional to $y$.
@@
TCitH. The general zero-knowledge proof of knowledge obtained from the TCitH framework is provided in Figure˜1. We give a brief overview below; for further details, see [8, 6]. The prover attempts to convince the verifier that they know a witness $\mathbf{w}\in\mathbb{F}_{p}^{k}$ satisfying $F\left(\mathbf{w}\right)=\mathbf{0}\in\mathbb{F}_{p}^{n}$ for a One-Way Function (OWF) $F$, described by the system of polynomial relations $F=(f_{1},\ldots,f_{m})\in\big(\mathbb{F}_{p}\left[x_{1},\ldots,x_{k}\right]_{\leq d}\big)^{m}$. To this end, the prover uses a GGM-based Vector Commitment (VC) scheme [17, 10] to sample vectors of random polynomials
$P_{u}\in\big(\mathbb{K}\left[X\right]_{\leq d-1}\big)^{\eta}\quad\text{and}\quad P_{w}\in\big(\mathbb{K}\left[X\right]_{\leq 1}\big)^{k}$
@@
with $P_{w}\left(\infty\right)\in\mathbb{F}_{p}^{k}$. The VC also generates a commitment $\mathsf{com}$ to $P_{w},P_{u}$ and the opening key $\mathsf{key}$. To embed the witness $\mathbf{w}$ into $P_{w}$, the prover computes the offset $\boldsymbol{\Delta}_{\mathbf{w}}=\mathbf{w}-P_{w}\left(\infty\right)\in\mathbb{F}_{p}^{k}$, for which $\tilde{P}_{w}\left(X\right)=P_{w}\left(X\right)+\boldsymbol{\Delta}_{\mathbf{w}}\cdot X$ has leading [^2] The witness can also be embedded into the constant coefficient; see [6]. coefficients $\mathbf{w}$. The offset $\boldsymbol{\Delta}_{\mathbf{w}}$ and commitment $\mathsf{com}$ are then sent to the verifier. Next, the polynomial relation $F$ is evaluated at $\tilde{P}_{w}\left(X\right)$. Specifically, for $i\in[n]$ and $\ell\leq d$, denote by $f_{i,\ell}(\mathbf{x})$ the degree-$\ell$ homogeneous component (i.e. the sum of all the monomials of degree $\ell$) of $f_{i}(\mathbf{x})$ . Then, $F(\tilde{P}_{w})(X)$ is defined as
$\Big(\sum_{\ell}X^{d-\ell}f_{1,\ell}(\tilde{P}_{w}\left(X\right)),\ldots,\sum_{\ell}X^{d-\ell}f_{n,\ell}(\tilde{P}_{w}\left(X\right))\Big).$
@@
Thus, $F(\tilde{P}_{w})(\infty)=F\left(\mathbf{w}\right)$, i.e., the embedded witness satisfies the polynomial relation $F$ if and only if $F(\tilde{P}_{w})(X)$ has degree at most d1d-1. Hence, $F(\tilde{P}_{w})(X)$ consists of $n$ polynomials of degree d1d-1. The number of polynomials is reduced through batching. To this end, the verifier samples a random matrix $\boldsymbol{\Gamma}\in\mathbb{K}^{\eta\times n}$ and the prover computes $\boldsymbol{\Gamma}\cdot F(\tilde{P}_{w})$, thereby reducing the number of polynomials from $n$ to $\eta$ while preserving the degree with overwhelming probability. The prover masks $\boldsymbol{\Gamma}\cdot F(\tilde{P}_{w})(X)$ by computing
$Q(X)=P_{u}\left(X\right)+\boldsymbol{\Gamma}\cdot F(\tilde{P}_{w})(X),\vskip-2.84526pt$
@@
where $P_{u}\left(X\right)$ is the length-$\eta$ vector of random polynomials of degree d1d-1 generated by the VC scheme. Sending the masked polynomial vector $Q(X)$ to the verifier does not reveal any information about the witness. To verify correctness, the verifier challenges the prover at a random evaluation point $r$ in the evaluation domain $\Omega\subseteq\mathbb{K}$ of size $N=\left\lvert\Omega\right\rvert$. The prover produces the corresponding opening $\mathsf{opening}_{r}$ using its opening key $\mathsf{key}$. The verifier checks $\mathsf{opening}_{r}$ against $\mathsf{com}$ and, in case of success, computes the evaluations $P_{w}\left(r\right)\in\mathbb{K}^{\eta}$ and $P_{u}\left(r\right)\in\mathbb{K}^{k}$. Then, the verifier computes $P_{u}\left(r\right)+\boldsymbol{\Gamma}\cdot F(\tilde{P}_{w}\left(r\right))$ and checks whether it matches $Q\left(r\right)$, where $Q(X)$ is the polynomial vector received from the prover.
Remark 1.
@@
In [6], also a three-pass variant of Figure˜1 is given, which is obtained by skipping the batching step. This simplifies the protocol, but increases the required communication, which is why we omit this variant.
@@
TCitH signature scheme. The proof system above has a soundness error given by
@@
Property 1.
For grinding parameter $w$, let the TCitH proof system parameters satisfy
@@
Then the resulting signature scheme is EUF-CMA-secure [6].
Signatures are of the form $\sigma=(\sigma_{\mathrm{sym}},\sigma_{\mathbf{w}},\sigma_{F})$, where $\sigma_{\mathbf{w}}$ corresponds to the witness, $\sigma_{F}$ to the OWF $F$, and $\sigma_{\mathrm{sym}}$ to the VC scheme. The components’ sizes [^3] In general, the sizes depend on the specific instantiation of the VC scheme [18], and optimizations not described here can be applied. The formulas given follow [6], to which we refer the reader for further details. are given by
$\left\lvert\sigma_{\mathrm{sym}}\right\rvert=\tau\cdot\lambda\cdot(\log_{2}(N)+1)+5\lambda+32,$
$\left\lvert\sigma_{\mathbf{w}}\right\rvert=\tau\cdot k\cdot\log_{2}(p),\text{ }\left\lvert\sigma_{F}\right\rvert=\tau\cdot\eta\cdot(d-1)\cdot\mu\cdot\log_{2}(p).\\$
VOLEitH. Instead of performing $\tau$ parallel repetitions, VOLEitH employs a technique that allows sampling $r$ from an evaluation domain $\Omega$ of size
@@
Avoiding parallel repetitions reduces the soundness error, which in turn enables shorter signatures. However, this comes at the cost of an additional round of interaction between the prover and the verifier to do a a consistency check. For further details, we refer to [5, 10].
@@
$\frac{N^{\tau}}{d}\geq 2^{\lambda-w}\quad\text{and}\quad p^{\rho}\geq 2^{\lambda},$
$\text{with}\quad\rho=\tau\cdot\left\lceil\log_{p}(N)\right\rceil,\quad B=\left\lceil\frac{16}{\log_{2}(p)}\right\rceil.$
@@
$\left\lvert\sigma_{\mathrm{sym}}\right\rvert$ $=\tau\cdot(\rho+B)\log_{2}(p)+\tau\cdot 2\lambda+T_{\mathrm{open}}\cdot\lambda+4\lambda+32,$
$\left\lvert\sigma_{\mathbf{w}}\right\rvert$ $=\tau\cdot k\cdot\log_{2}(p),$
$\left\lvert\sigma_{F}\right\rvert$ $=\tau\cdot(d-1)\cdot\rho\cdot\log_{2}(p).$
Similar to the TCitH framework, which requires $\eta\cdot\mu\cdot\log_{2}(p)\geq\lambda$, the VOLEitH framework requires $\rho\cdot\log_{2}(p)\geq\lambda$. However, VOLEitH allows for a smaller $\tau$, owing to its reduced soundness error, arguably at the cost of a more complex scheme.
TABLE I: Parameters of signatures based on $\mathcal{E}$-SDP obtained via the TCitH framework. The proof system parameters are selected analogously to those in MQOM [6] to ensure comparable runtime performance of the VC scheme.
Security category $\mathcal{E}$-SDP Proof System Performance